# Forecasting Electricity Market Signals

### v4

@@ 4.5 Area Control Error Forecasting for Reserve Market Participants
Methods NMSE NMAE MASE sMAPE
SNARX [4] (6) $1.1922$ (6) $1.1303$ (6) $0.7029$ (6) $0.4605$
WIAE-GPF (1) $\mathbf{0.5957}$ (1) $\mathbf{0.7555}$ (1) $\mathbf{0.4698}$ (1) $\mathbf{0.1059}$
TLAE [21] (5) $1.1727$ (5) $1.0605$ (5) $0.6595$ (3) $0.2782$
DeepVAR [11] (7) $1.4431$ (7) $1.1750$ (7) $0.7307$ (5) $0.3952$
BWGVT [19] (3) $0.9562$ (2) $0.9793$ (2) $0.6090$ (4) $0.3168$
Pyraformer [34] (4) $0.9783$ (4) $0.9948$ (4) $0.6186$ (7) $0.4986$
Informer [33] (2) $0.6006$ (3) $0.9819$ (3) $0.6106$ (2) $0.2247$
Table 8: Probabilistic Estimation Results of ACE forecasting for PJM. The prediction step is 5-minute ahead.
Methods CRPS CPE (90%) [NCW] CPE (50%) [NCW] CPE (10%) [NCW]
SNARX [4] (5) $2.1007$ (5) $-0.8260$ $[1.6759]$ (5) $-0.4797$ $[1.6803]$ (2) $0.0343$ $[5.5768]$
WIAE-GPF (1) $\mathbf{0.0081}$ (1) $\mathbf{-0.0016}$ $[0.9199]$ (1) $\mathbf{0.0321}$ $[0.9336]$ (1) $\mathbf{-0.0132}$ $[0.8885]$
TLAE [21] (4) $1.5541$ (4) $-0.7857$ $[0.0004]$ (4) $-0.4489$ $[0.0005]$ (4) $-0.0957$ $[0.0027]$
DeepVAR [11] (3) $1.2947$ (3) $-0.3526$ $[0.5665]$ (3) $-0.2560$ $[0.5296]$ (3) $-0.0521$ $[0.5434]$
BWGVT [19] (2) $1.2488$ (2) $0.0065$ $[1.8309]$ (2) $0.0754$ $[2.0385]$ (5) $0.0996$ $[2.4261]$
WIAE-GPF achieved better performance than other methods, with CRPS less than $0.01$ and sMAPE less than $11\%$, as shown by the WIAE-GPF row. We credited the strong performance of WIAE-GPF to the simplicity of its latent process, and its Bayesian sufficiency. BWGVT ranked second among all methods since the ACE data has few outliers. But its CRPS at $1.2488$ is dramatically larger than that of WIAE-GPF. Its CPE and NCW for 10% confidence interval prediction also showed that it cannot accurately predict a narrow interval. Pyraformer and Informer, trained with NMSE, had better performance under NMSE but worse under NMAE. With NMSE over $110\%$, TLAE had the worst performance among GPF methods. DeepVAR and SNARX performed worse than the other forecasting methods, with NMSE and NMAE larger than $110\%$, possibly due to model mismatch.
4.6 Discussion: On using LLM
@@
As LLM-based forecasting techniques, Pyraformer, Informer and BWGVT performed better than the more conventional SNARX (see Table 3, 5, 7). Compared with the more straightforward deep learning method of DeepVAR, LLM-based forecasting did not show clear advantages. The same can be said when they were compared with TLAE. Authors of [20] pointed out that the simple convolutional neural network outperformed RNNs and LLMs on imbalance price forecasting, for the forecasted time series are not a good fit to the complicated deep learning models.
@@ 5 Conclusion
This paper presents WIAE-GPF, a generative AI approach to probabilistic forecasting of nonparametric time series based on the innovation representation pioneered by Wiener, Kallianpur, and Rosenblatt six decades earlier. Three take-away conclusions stand out. One is that the innovation representation ensures that WIAE-GPF produces the correct conditional probability distributions under perfect learning condition assumptions. To our knowledge, WIAE-GPF is the first nonparametric GPF technique with such a theoretical guarantee. Second, WIAE-GPF demonstrated superior performance against major machine learning-based probabilistic forecasting techniques in our numerical studies using actual market data, including some of the advanced approaches involving transformer architecture, attention mechanism, and large language models. In addition, the local stationarity hypothesis assumed by the weak innovation representation appeared to hold well. Third, with Bayesian sufficiency established in this paper, Rosenblatt’s weak innovation representation of time series can be considered as a canonical representation and a powerful tool for stochastic decision making. Its applications in anomaly detection in power systems can be found in [47, 48, 49].
Finally, The black-box style of deep learning is often criticized for its lack of interpretability; some consider such methods as black magic that produces miraculous results. It is worth mentioning that WIAE-GPF has a highly intuitive and interpretable architecture parallel to that of the classic Kalman filtering. In particular, Kalman filtering extracts innovation representation as part of the measurement update, followed by the time-updated prediction according to the state-space model. WIAE-GPF extracts innovations by the weak innovation encoder and produces time-updated predictions through the weak innovation decoder. In this context, WIAE-GPF is a generalization of Kalman filtering to nonparametric and non-Gaussian settings.
Comments on the limitation and future work are in order. WIAE-GPF is derived based on the innovation representation of stationary processes. Extensions to certain classes of nonstationary processes would be the natural next step. Note that an innovation representation exists for nonstationary Gaussian processes with a time-varying state space model. Extension of WIAE-GPF to nonstationary time series under regime-switching models is also a natural next step, given the evidence of effective applications of regime-switching techniques in price forecasting. See [20, 19] and references therein.

### v5

@@ 4.4 Area Control Error Forecasting for Reserve Market Participants
Shown by Table. 6, WIAE-GPF achieved better performance than other methods, with CRPS less than $0.01$ and sMAPE less than $11\%$, as shown by the WIAE-GPF row. We credited the strong performance of WIAE-GPF to the simplicity of its latent process, and its Bayesian sufficiency. BWGVT ranked second among all methods since the ACE data has few outliers. But its CRPS at $1.2488$ is dramatically larger than that of WIAE-GPF. Its CPE and NCW for 10% confidence interval prediction also showed that it cannot accurately predict a narrow interval. Pyraformer and Informer, trained with NMSE, had better performance under NMSE but worse under NMAE. With NMSE over $110\%$, TLAE had the worst performance among GPF methods. DeepVAR and SNARX performed worse than the other forecasting methods, with NMSE and NMAE larger than $110\%$, possibly due to model mismatch.
4.5 Discussion: On using LLM for price forecasting
@@
As LLM-based forecasting techniques, Pyraformer, Informer and BWGVT somtimes performed better than TLAE (see Table 4). Compared with the more straightforward deep learning method of DeepVAR, LLM-based forecasting did not show clear advantages. Authors of [40] pointed out that the simple convolutional neural network outperformed RNNs and LLMs on imbalance price forecasting, for the forecasted time series are not a good fit to the complicated deep learning models.
@@ 5 Conclusion
This paper introduces WIAE-GPF, a generative AI method for probabilistic forecasting of nonparametric time series, building on the innovation representation developed by Wiener, Kallianpur, and Rosenblatt over six decades ago. Three key findings emerge from our research. First, the innovation representation enables WIAE-GPF to produce accurate conditional probability distributions, provided perfect learning conditions are met. To the best of our knowledge, WIAE-GPF is the first nonparametric generative probabilistic forecasting (GPF) technique with such theoretical guarantees.
Second, WIAE-GPF outperformed leading machine learning-based probabilistic forecasting methods in our numerical experiments using real-world market data. This includes advanced models employing transformer architectures, attention mechanisms, and large language models. Additionally, the local stationarity hypothesis, integral to the weak innovation representation, held robustly in our studies.
Third, this paper establishes Bayesian sufficiency for Rosenblatt’s weak innovation representation, validating it as a canonical framework for time series and a powerful tool for stochastic decision-making. Its potential applications in power system anomaly detection are explored in [43, 44, 45].
Finally, deep learning methods are often criticized for their black-box nature, with some viewing them as mysterious techniques that produce impressive yet opaque results. In contrast, WIAE-GPF offers a highly intuitive and interpretable architecture, closely paralleling the classic Kalman filter. Specifically, in Kalman filtering, the innovation is extracted during the measurement update, followed by time-updated predictions based on a state-space model. Similarly, WIAE-GPF extracts innovations using its weak innovation encoder and generates time-updated predictions via the weak innovation decoder. In this sense, WIAE-GPF can be seen as a generalization of Kalman filtering to nonparametric and non-Gaussian settings.
It is important to address the limitations and future directions of this work. WIAE-GPF is derived from the innovation representation of stationary processes, so extending it to certain classes of nonstationary processes would be a natural next step. Notably, an innovation representation exists for nonstationary Gaussian processes using time-varying state-space models. Expanding WIAE-GPF to handle nonstationary time series under regime-switching models is also a promising avenue, particularly given the demonstrated effectiveness of regime-switching techniques in price forecasting. See [40, 28] for relevant applications.